According to eMarketer, adults in the United States spent an average of 92 minutes per day on social media platforms FacebookThe answers are in red. Please answer all. I have a hard time understanding when to use the table
Answer #1

a) P(X < 68)

= r( - - 68 -

68 - 92 =PZ<a 29

= P(Z < -0.83)

= 0.2033

b) P(X > 65)

- تا - )r =

= P(Z > 65 - 92 = 29

= P(Z > -0.93)

=1-P(Z < -0.93)

=1-0.1762

=0. 38

c) P(85 < X < 110)

- P(55–4_-104)

29 29

= P(-0.24 <Z < 0.62)

= P(Z <0.62) - P(Z < -0.24)

= 0.7324-0.4052

= 0.322

d) P(X < x) = 0.65

201,p(*=4<==4) = 0.65

cor,P(2<5>) -066

or 292 or, -= 0.39 29

or, 2 = 0.39 * 29 + 92

or, r = 103